Hollydale's is a clothing store in East Park. It paid an annual dividend of $2.70 last year to its shareholders and plans to increase the dividend annually at 3.5​%. It has 430,000 shares outstanding. The shares currently sell for $19.39 per share. ​ Hollydale's has 12,000 semiannual bonds outstanding with a coupon rate of 8.5​%, a maturity of 22 years, and a par value of $1,000. The bonds are currently selling for $1,167.12 per bond. What is the adjusted WACC for​ Hollydale's if the corporate tax rate is 15%?

What is the adjusted WACC for​ Hollydale's if the corporate tax rate is 15​%?
Answer #1
Value Weight Cost
Equity     8,337,700 37.3% 17.9%
Debt 14,005,440 62.7% 7.00%
Total 22,343,140 WACC 10.41%

Value = No. of shares x Price

Weight = Value / Total

Cost of equity, re = D0 x (1 + g) / P + g = 2.70 x (1 + 3.5%) / 19.39 + 3.5% = 17.91%

Cost of debt, can be calculated using I/Y function on a calculator

N = 22 x 2 = 44, PV = -1,167.12, FV = 1,000, PMT = 8.5% x 1000 / 2 = 42.5

=> Compute I/Y = 3.5% (semi-annual)

=> Annual cost of debt, rd = 3.5% x 2 = 7%

WACC = wd x rd x (1 - tax) + we x re

= 62.7% x 7% x (1 - 15%) + 37.3% x 17.91%

= 10.41%